The vice president of operations of Recycling Industries is evaluating the performance of two divisions organized as investment centers. Invested assets and condensed income statement data for the past year for each division are as follows:

Business Division

Consumer Division

Sales

$2,200,000

$2,490,000

Cost of goods sold

1,260,000

1,335,000

Operating expenses

698,000

831,300

Invested assets

1,100,000

2,490,000

1. Prepare condensed divisional income statements for the year ended December 31, 20Y8, assuming that there were no service department charges.

Recycling Industries

Divisional Income Statements

For the Year Ended December 31, 20Y8

1

  

Business Division

Consumer Division

2

Sales

3

Cost of goods sold

4

Gross profit

5

Operating expenses

6

Income from operations

2. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investment for each division. If required, round your answers to one decimal place.

Profit Margin

Investment Turnover

ROI

Business Division

%

%

Consumer Division

%

%

3. If management desires a minimum acceptable return on investment of 18.00%, determine the residual income for each division. Use the minus sign to indicate a negative income. Round your answers to the nearest dollar.

Residual Income

Business Division

$

Consumer Division

$

4. Discuss the evaluation of the two divisions, using the performance measures determined in parts (1), (2), and (3).

On the basis of income from operations, the Division is more profitable. However, income from operations does not consider the amount of invested assets in each division. On the basis of the rate of return on investment, the Division is more profitable. Even though the Division has a higher profit margin, the Division has a higher investment turnover, which generates the higher rate of return on investment. On the basis of residual income, the Division is the more profitable of the two divisions.
Answer #1

1.

Recycling Industries

Divisional Income Statements

Particulars Business Division Consumer Division
Sales $                 2,200,000 $                 2,490,000
Less: Cost of goods sold $                 1,260,000 $                 1,335,000
Gross profit $                     940,000 $                 1,155,000
Less:Operating expenses $                     698,000 $                     831,300
Income from operations $                     242,000 $                     323,700

2.

2) Particulars Business Division Consumer Division
Income from operations (a) $                     242,000 $                           323,700
Sales (b) $                 2,200,000 $                       2,490,000
Invested assets (c ) $                 1,100,000 $                       2,490,000
Profit Margin d = a/b*100 11.00% 13.00%
Investment Turnover e =b/c 2 1
ROI = f = d*e 22.00% 13.00%

For the Year Ended December 31, 20Y8

3. Residual Income = Income from operations - Minimum acceptable rate of return * Asset Invested

Particulars Business Division Consumer Division
Income from operations (a) $                     242,000 $                           323,700
Invested assets (b ) $                 1,100,000 $                       2,490,000
Minimum acceptable return on investment (c ) 18% 18%
Minimum acceptable Income (d =b*c) $                     198,000 $                           448,200
Residual Income = e = a-d $                       44,000 $                         (124,500)
4) On the basis of residual income, Business Division is the more profitable of the two divisions.
On the basis of income from operations, Consumer Division is the more profitable of the two divisions.
On the basis of the rate of return on investment, the   Business Division is more profitable.